Fermi National Laboratory


Accelerator Update for June 27, 2005 - June 29, 2005

Monday June 27
The day shift began with Operations monitoring store 4235 and a small Recycler stash, with the Antiproton Source (Pbar) stacking, and with the following experiments taking beam: MiniBooNE, NuMI, and Meson MCenter.

The Meson MTest experiment resumed taking beam at 8:24 AM.

Operations prepared the Recycler for shot setup at 1:17 PM to stash antiprotons.

At 2:22 PM, a Linac expert began switching the Preaccelerator from the I- Source to the H- Source. He completed the switch by 2:26 PM.

The Recycler stash ended and Pbar resumed stacking around 2:40 PM.

At 4:31 PM, the NuMI beam permit tripped off when a power supply (E:HV101) tripped off.

Operators couldn’t reset it and called an expert.

At 5:45 PM, a MI-8 transfer line quadrupole power supply (I:Q847) tripped off. This was only the first of many times it tripped off during the shift. It reset each time, eventually. Operators reported that the building was hot. The Duty Mechanic investigated and reset the AC, but he noted that its filters were clogged and needed cleaning. These trips shut off beam to MiniBooNE.

The power supply expert found and repaired a failed capacitor for E:HV101. NuMI resumed taking beam at 6:28 PM.

A Pbar expert began a stacktail study at 7:07 PM.

A Booster expert began a damper study at 7:59 PM.

Tuesday June 28
The midnight shift began with Operations monitoring store 4235 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

MTest experimenters said that they were done with beam for the night at 2:08 AM.

At 2:10 AM, Operations reported that Pbar emittances were high. Operators worked on cooling, RF, and the tune and had the emittances back to normal by 3 AM.

Operations terminated store 4235 at 5:01 AM and then prepared D-Zero (D0) for an access. They came out of access at 6:15 AM.

Operations put the Tevatron (TeV) into a dry squeeze at 6:54 AM.

Operations put the TeV into shot setup at 7:27 AM.

At 7:44 AM, the TeV aborted its protons. Operators investigated, but couldn’t find a problem. At 7:48 AM, the TeV aborted again. Operators reported that this was happening on coalesced reverse injections.

Operations reported at 9:41 AM that they had loaded the TeV with protons. They completed loading the antiprotons at 10:03 AM, but a minute later the TeV suffered a sector B2 quench.

Pbar resumed stacking and MiniBooNE and NuMI resumed taking beam at 10:50 AM.

MTest and MCenter resumed taking beam at 11:14 AM.

Operators prepared the TeV for access at 12:08 PM for maintenance work and to allow the Collider Detector at Fermilab (CDF) and D0 experimenters to access their collision halls.

CDF and D0 came out of access at 2:06 PM. Operations began the TeV turn on.

At 2:45 PM, a Linac RF station (KRF1) tripped off for the first of four times in fifteen minutes. An operator contacted an expert.

Operations put the TeV through a dry squeeze at 4:02 PM.

TeV experts began a study period at 4:45 PM. At 7:28 PM, E:Q874 began tripping its breaker again. Operators investigated and found the service building very warm. Operators called the Duty Mechanic to reset the air conditioning.

The TeV experts completed their study at 8:56 PM.

Operations put the TeV into shot setup at 10:17 PM.

Wednesday June 29
The midnight shift began with the TeV in shot setup, and with MiniBooNE and NuMI taking beam.

At 1:16 AM, a Booster RF station (BRF12) tripped off on a ground fault.

Operations established store 4241 at 1:58 AM with an initial luminosity of 72.16E30. Pbar resumed stacking.

At 3:49 AM, Operations reported that BRF13 had been tripping off. It tripped off twice more and then operators decided to leave it off for a rest.

Operations prepared the Recycler at 6:33 AM to stash antiprotons.

The Plan for Wednesday
The plans for today are to keep store 4241 in until late this evening or early on the midnight shift and to continue stacking and stashing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/29/2005   email Fermilab